NPP Projection Close 1/29

To do this evening/ was completed:

Presentation Issues: Noticed during that last round in theater 1 (Send Help) the picture seemed a bit darker than usual. I thought it might be the "look" of the film since this particular moment was set at night but during one of the more brightly lit scenes I saw a slight the lamp flicker ever so slightly in the brightest part of the frame. Also of note the picture in theater 2 seems to have a magenta tint that is most noticeable during bright scenes, especially over pure white elements like text or credits. Neither thing seems like a hindrance at the moment but it would be worth continuing to keep an eye on over the weekend.

Preshow Clips/Playlist:

Trailer Updates:

Special Events/ AV set up: 9 to 5 tech went fine. Each song needed a volume boost so new cues were added and the playlist saved and synced. There was some confusion from MOD about the pre-show length but that was cleared up quickly (in short the cue sheets did not specify the performance after the intro hence the question regarding the duration of the pre-show). I noticed trailers in the playlist and cut them out. The event memo requested no trailers. We opened the house roughly 7 minutes behind schedule and I got the show up to speed from the booth. As for the actual run of show I was asked to raise volume levels again on every song and in turn ended up raising the volume for everything else just to be safe (ex: the curtain call and feature...moving forward we should just build out the pre-show for these events with louder than necessary volumes knowing the audience is going to be so raucous and the levels will change regardless of the techs). Performer marks and light cues were all hit with aplomb. It was a rowdy crowd as usual and the film played without any issues. 

Ordering/Receiving (posters & trailers): Arco and How to Make a Killing posters were delivered and stored behind 3/7

Ordering/Receiving (content):

Outgoing/Shipping (content):

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning: Placed all equipment used for the event in theater 2 back beneath the curtain after the show

Cue sheets updated/checked (every shift): Yes

Marquee updated?: Added letters for Arco (opening tomorrow) and Send Help

Posters updated?: Placed a "now showing" snipe above Send Help outside

Cleaning lights on for the night?: Yes

Projection cart by entrance?: Yes

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will

NPP Projection Admin 7/6